# Auditory–Tactile Congruence for Synthesis of Adaptive Pain Expressions in RoboPatients

### v1

@@ 3.2 Abdominal Phantom
A force sensor platform and silicon phantom replicated a human’s abdomen as shown in Figure 1. The force sensor platform was developed to measure the palpation force during user palpation. The force sensor platform was made of four single-axis (20kg) load cells (Haljia, Inc.). The load cells were mounted at the four corners of a generic square shape (250 $\times$ 250 mm) acrylic platform. The distances between two adjacent load cells were 246 mm. A National Instruments DAQ (USB-6211) with MATLAB was used to acquire the signals from the load cells at a sampling rate of 1000 Hz. The sum of the force values of each load cell was used to calculate the total palpation force ($F_{tp}$). To minimize noise in the total palpation force, values below 1N were neglected based on observations from pilot experiments.
@@
where $F_{j}$ is the force reading from the $jth$ load cell mounted on the four corners of the force sensor platform.
A square shape (200 $\times$ 200 $\times$ 30 mm) silicone block was fabricated with Ecoflex 00-30 (Smooth-On, Inc.). We placed the silicone block on the force sensor platform. It was used as the abdominal phantom of the simulated human.
@@ 3.3 MorphFace
MorphFace is a controllable 3D physical-virtual hybrid face that can represent facial expressions of pain displayed on six different ethnicity-gender face identities: female and male faces of White, Black, and Asian ethnicity. The pain facial expressions are realized by 4 facial action units: AU4 (Brow Lowerer, AU7 (Lid Tightener), AU9 (Nose Wrinkler), and AU10 (Upper Lip Raiser), with the pain intensity $(PI)$ ranging from 0 (no pain) to 100 (maximum pain). More details of MorphFace can be found in our recent work. We included only the white male and white female face identities for this study. The servo motors attached to different parts of the face used to realise action units, were not actuated for this study. Hence the physical face was static while the projected facial expression was dynamic over time, based on the pain intensity.
@@ 4 Experiments
We asked 20 healthy participants (10 males and 10 females) aged between 21 and 50 years (mean age 30.05 years, SD 9.5 years). Students and faculty with no prior experience in pain to palpation mapping and no history of sensorimotor impairments were chosen to participate in this experiment. Refer to Figure 2A for the experiment setup.
The experiment consists of one task. The participant must palpate the abdominal phantom until they reach a target force. Four levels of target forces were presented to each participant to ensure they applied diverse forces when palpating. A pain sound was played when the applied force was equal to or greater than the target force. The participant must then rate whether they ‘strongly agree’ (a), ‘agree’ (b), ‘strongly disagree’ (c), or ‘disagree’ (d) through the selection interface. We labeled each choice with a letter (a-d) to make data management more intuitive.
For the experiments, four parameters were changed between the trials. The target force, pain sound, pain sound amplitude, and pain sound pitch. Four different values were used for the target forces, pain sound amplitude, and pain sound pitch. We determined these values would be appropriate from pilot trials.
@@
192 trials were conducted for each gender ( male and female). The total number of trials per participant was 384. The experiment was divided into two parts: male and female. In each part, the faces displayed on MorphFace and pain sounds were used according to gender. A matrix of all possible combinations of amplitude, pitch, force levels, and pain sounds was created in MATLAB. A combination of the four parameters was selected once at random from the matrix for each trial. None of the combinations were repeated for that gender.
Each trial lasted 3 seconds, and initialising the next trial took 1 second. There was no time limit for the participant to make their decision. During the 3-second trial, the total palpation force was acquired through the DAQ. When a trial began, the participant was presented with a graphical interface with a green colored marker (target force) and a face image on MorphFace. A dynamic green bar appeared and changed length when the participant applied force to the silicone block. Once the participant reached the target force, the pain sound was played through headphones (Razer Kraken X Lite), and the dynamic green bar became static and changed to red. The program then waited till the participant rated the pain sound. The program then moved on to conduct the next trial. Every 64 trials, the participant was given a 90-second break. At the end of the 192 trials, a 5-minute break was given before starting the trials for the next gender.
@@
The participants were allowed to familiarize themselves with the system. To do this, six familiarization trials were conducted per gender. None of the parameters were randomized during these trials. The pain sounds were played without alteration. The force levels changed in ascending order for each trial. The force levels for the 5th and 6th trials were 5 N and 10 N. Counterbalancing was introduced to this study by changing the order in which the participants began. For example, if a participant first started the experiment with male pain expressions, the next participant started the experiment with female pain expressions. This was to reduce bias, and the experiment took approximately one hour to complete for each participant. This experimental process is shown in Figure 2B.
@@ 5.1 Dataset
The dataset consists of the decisions collected from 20 individuals who interacted with our autonomous system. Each participant’s data comprised multiple features capturing the force information from palpation, corresponding pain sound, amplitude, and pitch generated by the robot and participant’s choice (agree-disagree). A section of this dataset is shown in Table 1. In this section, we present the results of our analysis, including dimensionality reduction, clustering, association rule mining, and the interpretation of key findings from these tests.
@@ 5.2 Results
The study’s primary aim was to explore the relationship between the acoustic properties of sound and the applied palpation force. The median and mode amplitude and pitch values for each target force are presented in Figure 3A and B. This analysis was conducted for trials where the participants chose ‘strongly agree’ and ‘agree’. The observed results show that the preferred median amplitude of pain sounds is at a minimum at the beginning, and reaches a maximum after 10 N and remains constant beyond that. The preferred median pitch remains constant until 10 N and falls to a minimum beyond that. This can be observed with both male and female pain sounds. This may be due to several factors that contradict intuitive expectations:
@@
Similarly, for female pain sounds, we observe that the preferred mode amplitude of pain sounds is at a minimum at the beginning, and reaches a maximum after 10 N and remains constant beyond that. The preferred mode pitch gradually decreases until 15 N and remains at a minimum beyond that. However, a different trend can be observed for mode amplitude and pitch values for male pain sounds. Both amplitude and pitch start at a minium when the target force is 5 N. They reach a maximum at 10 N and fall when the target force is 15 N. Pitch remains constant beyond that, and amplitude increases to 1.
Hence, these findings suggest that subjective and perceptual factors, as well as the unique acoustic properties of male and female pain sounds, play a significant role in shaping participants’ preferences. Further research could investigate how sound properties are processed at higher force levels and whether these preferences are influenced by cultural, emotional, or physiological expectations of pain representation.
The response percentages for agree/strongly agree and disagree/strongly disagree across all trials are shown in Figure 4. From the given responses—agree/strongly agree: 50.9375% , and disagree/strongly disagree: 49.0625%, near equal opinions could be observed for both cases. Hence, the responses are nearly evenly split between agreement and disagreement. The close split suggests that the task or question may be subjective, leading to varied interpretations or individual differences in how participants perceive and match pain sounds to force levels. Furthermore, from these results, we assume that an adequate number of choices of pain sounds were covered during the experiment.
@@
Figure 4: The responses given by the participants (Agree/Strongly agree and Disagree/Strongly disagree) for all the trials throughout the experiment are presented as percentages.
When the actual force distribution was plotted against each target force level, a slight increase in actual force was observed as the target force levels increased. This trend is illustrated in Figure 5 A and B. Considering the peak force distribution and mean values of these two plots, this slight increment becomes apparent, with female participants exerting higher forces on average compared to male participants. The observed higher forces exerted by female participants suggest differences in force perception, motor control, or task approach between genders. These variations highlight the need to account for individual or group differences in force-related tasks, as perception and motor strategies may differ.
@@
Figure 5: Violin plot visualizing the distribution of actual force responses across different target forces. The width of each violin reflects the density of actual forces, with the central line indicating the median and the shaded area showing the interquartile range (IQR). This visual highlights differences in actual force distributions among the four target force categories. The plots represent the force responses for A) male and B) female participants, respectively.
@@
A similar trend was observed in each gender’s response to pain sounds of their own gender, as shown in Figure 6. However, when examining responses to pain sounds of the opposite gender, the scatter increased significantly, with the data spread over a wider range in the violin plot. This may be attributed to differences in familiarity or emotional connection, as participants might find it harder to relate to or interpret the pain sounds of the opposite gender. Additionally, cognitive biases or emotional factors could influence how pain sounds are processed, leading to less consistent responses. These findings underscore the importance of accounting for gender-based perceptual differences in studies involving pain sounds, as they can significantly affect the interpretation of results and the design of such experiments.
@@
Figure 6: Violin plot visualising the distribution of actual force responses across different target forces. The width of each violin reflects the density of actual forces, with the central line indicating the median and the shaded area showing the interquartile range. This visual highlights differences in actual force distributions among the four target force categories. The plots represent the force responses of A) male participants on female pain sounds, B) male participants on male pain sounds, C) female participants on female pain sounds, and D) female participants on male pain sounds, respectively.
@@
Interestingly, a similarly scattered pattern was observed in the overall responses to male pain sounds, as shown in Figure 7. Despite the variability, the distribution peaks consistently followed the same trend: a subtle increase in actual force with rising target force levels. The scattered response pattern for male pain sounds, combined with consistent peaks showing a subtle increase in actual force with higher target levels, suggests underlying consistency amid variability. Despite individual differences, the predictable trend indicates a shared perception of the relationship between force and pain sounds. This highlights the reliability of the task design in eliciting responses aligned with target force levels, even when sound characteristics introduce variability.
@@
Figure 7: Violin plot visualising the distribution of actual force responses across different target forces. The width of each violin reflects the density of actual forces, with the central line indicating the median and the shaded area showing the interquartile range. This visual highlights differences in actual force distributions among the four target force categories. The plots represent the force responses of all participants, despite gender, on A) female pain sounds and B) male pain sounds, respectively.
@@
Figure 8: A) Frequency of participants’ choices plotted against force thresholds. B) PCA of auditory and force features. Both plots are colored by the participant’s choice.
@@ 5.3 Further analysis upon human feedback
Figure 8A illustrates the distribution of force threshold values across participants’ choices. It can be observed from the figure that Choice ‘b’ (agree) is the most frequent response across all force levels, and Choice ‘a’ (strongly agree) becomes more prominent at higher thresholds. The most likely reason for this behavior was increased perceived confidence or certainty by participants as the force threshold increases. At higher force levels, stimuli may become more easily perceptible or unambiguous, leading participants to respond with greater agreement or stronger affirmation (i.e., choosing ‘a’ – strongly agree) instead of just ‘b’ – agree. This aligns with typical psychophysical response patterns where stronger stimuli yield more confident responses.
@@
We then performed one-way ANOVA tests to assess whether the amplitude, pitch, force threshold, and pain sound varied significantly across participants’ choices. Table 2 summarizes these results. This analysis revealed significant effects for amplitude (F = 29.65, p $<$ 0.001), pitch (F = 3.92, p = 0.008), and force threshold (F = 30.24, p $<$ 0.001), while pain sound did not differ significantly across participant responses (F = 0.44, p = 0.722). Post-hoc Tukey HSD tests further clarified these effects: for amplitude, significant differences were observed between ‘b’ (agree) and both ‘c’ (disagree) and ‘d’ (strongly disagree), as well as between ‘d’ and ‘a’ (strongly agree). This indicates that louder sounds were more likely to elicit agreement, while quieter or mismatched amplitudes led to disagreement. In the case of pitch, the significant contrasts were between ‘b’ and ‘c’, and between ‘c’ and ‘d’, suggesting that pitch plays a more nuanced role in participant perception. For force thresholds, the most notable finding was that ‘a’ responses differed significantly from all other choices, implying that participants who strongly agreed tended to apply greater force. Overall, these results suggest that both auditory intensity and tactile input (force) strongly influenced perceived congruence, while speaker identity (pain sound) has little perceptual impact.
Table 3: Principal component loadings for the four key features: amplitude, pitch, pain sound and force threshold.
Feature PC1 PC2
Amplitude 0.73 0.05
Pitch -0.44 -0.31
Pain Sound 0.50 -0.61
Force Threshold -0.18 -0.73
Table 4: Contingency table showing the distribution of participant choices across clusters.
Cluster ‘a’ ‘b’ ‘c’ ‘d’ Total
0 441 1281 593 1077 3392
1 461 1796 581 1450 4288
Total 902 3077 1174 2527 7680
To further investigate the underlying structure of the feature space, we applied Principal Component Analysis (PCA) to the four variables: amplitude, pitch, pain sound and force threshold. The first two principal components together accounted for approximately $60\%$ of the total variance (PC1 = $34.6\%$, PC2 = $25.3\%$). Examination of component loadings shown in Table 3, revealed that PC1 was driven primarily by amplitude (positively) and pitch (negatively), while PC2 reflected a combination of force threshold and pain sound. These findings suggest that the feature space can be interpreted in terms of a loudness-pitch axis (PC1) and a force-sound identity axis (PC2).
@@
As shown in Figure 8B, participant responses form overlapping but non-random clusters in PCA space. Notably, data points associated with choice ‘b’ (agreement) appear more densely centered near the origin, suggesting that moderate values of pitch and amplitude are generally associated with perceptual agreement. In contrast, choice ‘a’ (strong agreement) and choice ‘d’ (strong disagreement) tend to appear at opposite poles along PC1, reinforcing the idea that loudness and pitch polarity modulate perceived sound-force congruence. Responses marked as choice ‘c’ (disagree) are more scattered but seem to occupy regions with lower PC1 and moderate-to-high PC2 values, possibly reflecting mismatches in perceived force or unnatural sound identity. This spatial distribution supports the hypothesis that combinations of sound intensity and pitch, more than sound identity alone, contribute to consistent perceptual judgments of auditory feedback in palpation scenarios.
To build on the structure revealed by PCA, we applied K-Means clustering to the standardized feature set consisting of amplitude, pitch, pain sound, and force threshold. Using silhouette score analysis, we found that a two-cluster solution (k = 2) provided the optimal balance between cohesion and separation. As visualized in PCA space in Figure 9A, the resulting clusters occupy partially overlapping but distinguishable regions. Cluster 0 (in blue) is more concentrated in the positive PC1 region—previously associated with higher amplitude and lower pitch—while Cluster 1 (in orange) dominates the left side of the plot, corresponding to lower amplitude and potentially pitch-mismatched combinations. This spatial separation reflects meaningful perceptual groupings embedded in the sound-force feature space.
To test whether these data-driven clusters align with participant judgments, we compared cluster membership to choices using a chi-square test of independence. The contingency table (Table 4) revealed a significant non-random association between clusters and participant responses ($\chi^{2}$ = 37.80, p $<$ 0.001). Specifically, cluster 1 was more frequently associated with responses indicating stronger agreement or disagreement (‘b’ and ‘d’), while cluster 0 showed a greater proportion of moderate or less confident responses (‘a’ and ‘c’). This suggests that perceptual extremity in response patterns may correspond to more distinct auditory-force pairings in the underlying feature space.
To further investigate which features most strongly contributed to the observed clustering patterns, we trained a Random Forest classifier to predict cluster membership using the original input features: amplitude, pitch, pain sound and force threshold. The resulting importance scores, visualized in Figure 9B, clearly indicate that pitch is the most influential predictor, accounting for approximately $70\%$ of the decision-making weight. Amplitude followed as the second most important feature, contributing around $30\%$, while both force threshold and pain sound contributed negligibly to the model’s performance.
These findings reinforce the interpretation that auditory properties—particularly pitch and loudness—are central to how participants differentiate perceptual responses, aligning with earlier results from the ANOVA and PCA analyses. Despite its role as a tactile input, force threshold appears to play a less defining role in how perceptual clusters emerge from the data, and pain sound (speaker identity) again shows minimal influence. Together, this supports the conclusion that perceptual judgments of pain-related sounds in this context are shaped primarily by core acoustic features, rather than the physical input or voice characteristics. These insights offer practical implications for the design of robopatient auditory feedback systems, suggesting that careful modulation of pitch and amplitude may be more effective than adjusting speaker voice or mapping based on force input alone.
Figure 9: A) PCA plot with points colored by K-Means cluster (k=2) B) Random Forest feature importance for cluster prediction

### v2

@@ 3.2 Abdominal Phantom
A force sensor platform and silicon phantom replicated a human’s abdomen as shown in Figure 1. The force sensor platform was developed to measure the palpation force during user palpation. The force sensor platform was made of four single-axis (20 kg) load cells (Haljia, Inc.). The load cells were mounted at the four corners of a generic square shape (250 $\times$ 250 mm) acrylic platform. The distances between two adjacent load cells were 246 mm. A National Instruments DAQ (USB-6211) with MATLAB was used to acquire the signals from the load cells at a sampling rate of 1000 Hz. The force value of each load cell was used to calculate the total palpation force ($F_{tp}$). To minimize noise in the total palpation force, values below 1N were neglected based on observations from pilot experiments.
@@
where $F_{j}$ is the force reading from the jthjth load cell mounted on the four corners of the force sensor platform.
A square-shaped (200 $\times$ 200 $\times$ 30 mm) silicone block was fabricated with Ecoflex 00-30 (Smooth-On, Inc.). We placed this silicone block on the force sensor platform. It was used as the abdominal phantom of the simulated human.
@@ 3.3 MorphFace
MorphFace is a controllable 3D physical-virtual hybrid face that can represent facial expressions of pain displayed on six different ethnicity-gender face identities: female and male faces of White, Black, and Asian ethnicity. The pain facial expressions are realized by 4 facial action units: AU4 (Brow Lowerer, AU7 (Lid Tightener), AU9 (Nose Wrinkler), and AU10 (Upper Lip Raiser), with the pain intensity $(PI)$ ranging from 0 (no pain) to 100 (maximum pain). More details of MorphFace can be found in our recent work [25]. We included only the white male and white female face identities for this study. The servo motors attached to different parts of the face used to realise action units, were not actuated for this study. Hence the physical face was static while the projected facial expression was dynamic over time, based on the pain intensity.
@@ 4 Experiments
We asked 20 healthy participants (10 males and 10 females) aged between 21 and 50 years (mean age 30.05 years, SD 9.5 years) to take part in this study. Students and faculty with no prior experience in pain to palpation mapping and no history of sensorimotor impairments were chosen to participate in this experiment. Refer to Figure 2A for the experiment setup. We conducted our study with an equal number of participants from each gender to keep gender balance and study gender-related trends as well [44].
The experiment consists of one task. The participant must palpate the abdominal phantom until they reach a target force. Four levels of target forces were presented to each participant to ensure they applied a diverse range of forces when palpating. A pain sound was played when the applied force was equal to or greater than the target force. We refer to the set target level as the Force Threshold (5/10/15/20 N) and to the measured total palpation force as $F_{tp}$ (in N). The participant must then rate whether they ‘strongly agree’ (a), ‘agree’ (b), ‘strongly disagree’ (c), or ‘disagree’ (d) through the selection interface. We labeled each choice with a letter (a-d) to make data management more intuitive. We aimed to explore the relationship between palpation force and the creation of a palpation-to-pain sound mapping across a range of forces. Therefore, the target force was used as a trigger or control variable. In a previous attempt to map pain responses to palpation force [46], we observed complex relationships between physical palpation and auditory pain expressions that contradicted our initial hypothesis. Using proximal policy optimization (PPO), the study learned individual preferences that linked specific forces to pain expressions, indicating that personalized mappings were more effective than a generalized model. However, the study involved only a small number of participants, motivating the present expanded investigation into the relationship between physical and expressive aspects of pain.
For the experiments, four parameters were changed between the trials. The target force, pain sound, pain sound amplitude, and pain sound pitch. Four different values were used for the target forces, pain sound amplitude, and pain sound pitch. We determined these values would be appropriate from pilot trials. We only chose three individual pain sounds for this scoping study, as the combination of pitches and amplitudes with relation to an individual pain sound can be tremendously large and is very difficult to study. Furthermore, we only chose to change these four significant parameters during the trials, as changing no more than four variables simultaneously ensures it is within the human capacity to reliably interpret and respond during affective or perceptual tasks [4],[18].
@@
192 trials were conducted for each gender ( male and female). The total number of trials per participant was 384. The experiment was divided into two parts: robopatient displaying male and female expressions at a time. In each part, the facial expressions displayed on MorphFace and pain sounds played were selected based on a particular gender. This was to minimize cognitive load, which can be introduced by randomly changing the facial and vocal expressions of the robopatient to the opposite gender during the trials. A matrix of all possible combinations of amplitude, pitch, force levels, and pain sounds was created in MATLAB. A combination of the four parameters was selected once at random from the matrix for each trial. None of the combinations were repeated for that gender.
Each trial lasted 3 seconds, and initializing the next trial took 1 second. There was no time limit for the participants to make their decision. During the 3-second trial, the total palpation force was acquired through the DAQ. When a trial began, the participant was presented with a graphical interface with a green colored marker (target force) and a face image on MorphFace. A dynamic green bar appeared, adjusting its length in real time according to the force applied during palpation of the silicone block. Once the participant reached the target force, the pain sound was played through headphones (Razer Kraken X Lite), and the dynamic green bar became static and turned red. The program then waited for the participant to rate the pain sound. We asked the participants to choose their response based on how appropriate the pain sound they heard was in relation to the palpation force (i.e., strongly agree, agree, strongly disagree, and disagree). The program then moved on to the next trial. Every 64 trials, the participant was given a 90-second break. At the end of the 192 trials, a 5 minutes break was given before starting the trials for the next gender.
The participants were allowed to familiarize themselves with the system. To do this, six familiarization trials were conducted for each gender. None of the parameters were randomized during these trials. The pain sounds were played without alteration. The force levels changed in ascending order for each trial. The force levels for the 5th and 6th trials were 5 N and 10 N. Counterbalancing was introduced to this study by changing the order in which the participants began. For example, if a participant first started the experiment with male pain expressions, the next participant started the experiment with female pain expressions. This was to reduce bias, and the experiment took approximately one hour to complete for each participant. This experimental process is shown in Figure 2B.
The participants spent well over 37 minutes on the main experiments, which include the four 90-second breaks and the 5-minute break. They spent more than 48 seconds on the familiarization trials. The exact experiment time of each participant depends on their decision time. However, none of the participants took more than an hour to complete the study, which included their briefing.
@@
Though the main focus of this scoping study is exploring audio-tactile congruence, we believe the integration of facial expressions is necessary alongside the vocal pain sounds during the experiments. Facial pain expressions would provide context and help participants perceive the pain sounds effectively; otherwise, the sounds on their own can be odd and confusing. Despite the integration of facial pain expressions, we believe they would not affect the nature or findings of this study, as the responses of the participants would be dependent on the pain sounds.
@@ 5.1 Dataset
The data of each participant comprised multiple features, capturing the force information from palpation, corresponding pain sound, amplitude, and pitch generated by the robot and participant’s choice (agree-disagree). A section of this dataset is shown in Table 1. In this section, we present the results of our analysis, including dimensionality reduction, clustering, association rule mining, and the interpretation of key findings from these tests.
@@ 5.2 Results
The study’s primary aim was to explore the relationship between the acoustic properties of sound and the applied palpation force. The median and mode amplitude and pitch values for each target force are presented in Figure 3A and B. This analysis was conducted for trials where the participants chose ‘strongly agree’ and ‘agree’. The observed results show that the preferred median amplitude of male pain sounds is at a minimum at the beginning, and reaches a maximum after 10 N and remains constant beyond that. The preferred median pitch remains constant until 10 N and falls to a minimum beyond that. Similarly, for female pain sounds, we observe that the preferred mode amplitude of pain sounds is at a minimum at the beginning, and reaches a maximum after 10 N and remains constant beyond that. The preferred mode pitch gradually decreases until 15 N and remains at a minimum beyond that. However, a different trend can be observed for mode amplitude and pitch values for male pain sounds. Both amplitude and pitch start at a minimum when the target force is 5 N. They reach a maximum at 10 N and fall when the target force is 15 N. Pitch remains constant beyond that, and amplitude increases to 1. Hence, this trend can be observed with both male and female pain sounds.
These findings suggest that subjective and perceptual factors, as well as the unique acoustic properties of male and female pain sounds, play a significant role in shaping participants’ preferences. Further research could investigate how sound properties are processed at higher force levels and whether these preferences are influenced by cultural, emotional, or physiological expectations of pain representation.
However, intuitively, pain induced by an applied force would be expected to increase with both sound pitch and amplitude. Several factors may account for the discrepancy:
@@
These features have been observed in previous pain-related studies, but larger-scale experiments will be needed to confirm these behaviors in robotic patients.
To gain a more discrete understanding of whether participants agreed or disagreed with the pain sound combinations we used in the experiments, we combined the strongly agree/ agree and strongly disagree/ disagree to create two distinct groups. This is done according to the common response collapsing scheme [50], where we collapsed the four response options into two (agree/disagree) for binary analysis. The response percentages for agree/ strongly agree and disagree/ strongly disagree across all trials are shown in Figure 4A. From the given responses—agree/ strongly agree: 50.9375%, and disagree/strongly disagree: 49.0625%, near equal opinions could be observed for both cases. Hence, the responses are nearly evenly split between agreement and disagreement. The close split suggests that the task or question may be subjective, leading to varied interpretations or individual differences in how participants perceive and match pain sounds to force levels. This is more clearly seen in Figure 4B, where a large variation can be observed between the split of strongly agree and agree compared to strongly disagree and disagree. Furthermore, from these results, we assume that an adequate number of choices of pain sounds were covered during the experiment.
@@
Figure 4: A) The responses given by the participants (Agree/Strongly agree and Disagree/Strongly disagree) for all the trials throughout the experiment are presented as percentages. B) The response probability of each response is shown here.
When the actual force distribution was plotted against each target force level, a slight increase in actual force was observed as the target force levels increased. This trend is illustrated in Figures 5 A and B. Taking into account the peak force distribution and mean values, we observed a small increase in the forces exerted by the female participants. We saw this pattern in another study [46] that we conducted within the same context as well. Hence, this could potentially suggest differences in force perception, motor control, or task approach between genders. These variations may highlight the need to account for individual or group differences in force-related tasks, as perception and motor strategies may differ. A larger sample size of participants in future work may further reinforce these conclusions. A detailed analysis of gender-based differences in responding to diverse robotic pain expressions is provided in [46]. In brief, variations in initial responses can stem from physiological and psychosocial factors, consistent with prior work showing that gendered vocal cues shape perceived urgency, pain tolerance, and response behavior. The work in [46] further shows that adaptive expressions and real-time human feedback can address this issue.
@@
Figure 5: Violin plot visualizing the distribution of actual force responses across different target forces. The width of each violin reflects the density of actual forces, with the central line indicating the median and the shaded area showing the inter-quartile range (IQR). This visual highlights differences in actual force distributions among the four target force categories. The plots represent the force responses for male and female participants for each target force, respectively (refer to legend). The dashed black lines represent the mean actual force of each target force.
@@
A similar trend was observed in each gender’s response to pain sounds of their own gender, as shown in Figure 6A-D. However, when examining responses to pain sounds of the opposite gender, the scatter increased significantly, with the data spread over a wider range in the violin plot. This may be attributed to differences in familiarity or emotional connection, as participants might find it harder to relate to or interpret the pain sounds of the opposite gender. Additionally, cognitive biases or emotional factors could influence how pain sounds are processed, leading to less consistent responses. These findings underscore the importance of accounting for gender-based perceptual differences in studies involving pain sounds, as they can significantly affect the interpretation of results and the design of such experiments.
@@
Figure 6: Violin plot visualising the distribution of actual force responses across different target forces. The width of each violin reflects the density of actual forces, with the central line indicating the median and the shaded area showing the interquartile range. This visual highlights differences in actual force distributions among the four target force categories. The plots represent the force responses of A) male participants on female pain sounds, B) male participants on male pain sounds, C) female participants on female pain sounds, and D) female participants on male pain sounds, respectively. The dashed black lines represent the mean actual force of each target force.
@@
Interestingly, a similarly scattered pattern was observed in the overall responses to male pain sounds, as shown in Figure 7. Despite the variability, the distribution peaks consistently followed the same trend: a subtle increase in actual force with rising target force levels. The scattered response pattern for male pain sounds, combined with consistent peaks showing a subtle increase in actual force with higher target levels, suggests underlying consistency amid variability. Given the variability across the participants, our findings potentially indicate a relationship between applied force and perceived pain sounds; further analysis is required to confirm this. This potentially highlights the need for a task design that clearly elicits responses aligned with target force levels, even when sound characteristics introduce variability.
It should be noted that in Figures 5,6, and 7, participants apply forces that are higher and lower than the target force levels. We believe that participants can subjectively determine the necessary amount of force they would need to apply within the time limit to reach the target force. This could result in the maximum palpation force being higher or lower than the target force. Hence, some of the trials may result in the maximum palpation force being below the target force. While the actual forces differed from the target forces in many cases, we believe that the participants attempted to palpate in a way that matched the target force. Accordingly, as depicted in Figure 3, the results show a causal relationship in which amplitude increases with higher target forces, while pitch decreases. However, more trials with a larger number of participants can improve the insight into these findings.
@@
Figure 7: Violin plot visualising the distribution of actual force responses across different target forces. The width of each violin reflects the density of actual forces, with the central line indicating the median and the shaded area showing the interquartile range. This visual highlights differences in actual force distributions among the four target force categories. The plots represent the force responses of all participants, despite gender, on A) female pain sounds and B) male pain sounds, respectively. The dashed black lines represent the mean actual force of each target force.
@@
Figure 8: A) Frequency of participants’ choices plotted against force thresholds. B) Exploratory PCA (visualization only; PCs from z-scored features) of auditory and force features. Both plots are colored by the participant’s choice.
@@ 5.3 Further analysis upon human feedback
We analyze four predictors: Amplitude (unitless loudness scaling), Pitch (ratio levels), PainSound identity (categorical), and Force Threshold (the discrete target condition: 5/10/15/20 N). The total palpation force measured from the load cells is denoted $F_{tp}$ (in N). When principal components (PC1/PC2) are referenced, they are computed from standardized features (z-scores) and used for visualization only; all statistical inference is performed in the native feature space (Amplitude, Pitch, PainSound, Force Threshold). One-way ANOVA $F$ statistics are defined as $F=\mathrm{MS}_{\text{between}}/\mathrm{MS}_{\text{within}}$, and post-hoc pairwise contrasts use Tukey’s HSD (based on the studentized-range distribution).
@@
Figure 8A illustrates the distribution of force threshold values across participants’ choices. It can be observed from the figure that Choice ‘b’ (agree) is the most frequent response across all force levels, and Choice ‘a’ (strongly agree) becomes more prominent at higher thresholds. The most likely reason for this behavior was increased perceived confidence or certainty by participants as the force threshold increases. At higher force levels, stimuli may become more easily perceptible or unambiguous, which leads participants to respond with greater agreement or stronger affirmation (i.e., choosing ‘a’ – strongly agree) instead of just ‘b’ – agree. This is further reinforced by ‘d’ (strongly disagree) being the second frequently selected response for every force threshold. This consistent presence of strong disagreement suggests a subgroup of participants who perceived the stimuli as insufficient or ambiguous at lower thresholds, with frequency remaining stable from 5–15 N before decreasing at 20 N. Including both trends provides a broader picture of the response distribution and aligns with typical psychophysical patterns where stimulus strength influences both certainty and polarity of responses. This aligns with typical psychophysical response patterns where stronger stimuli yield more confident responses [52, 9, 31].
@@
We then performed one-way ANOVA tests to assess whether the amplitude, pitch, force threshold, and pain sound varied significantly across participant choices. Table 2 summarizes these results. This analysis revealed significant effects for amplitude (F = 29.65, p << 0.001), pitch (F = 3.92, p = 0.008), and force threshold (F = 30.24, p << 0.001), while pain sound did not differ significantly across participant responses (F = 0.44, p = 0.722). Post-hoc Tukey HSD tests further clarified these effects: for amplitude, significant differences were observed between ‘b’ (agree) and both ‘c’ (disagree) and ‘d’ (strongly disagree), as well as between ‘d’ and ‘a’ (strongly agree). This indicates that louder sounds were more likely to elicit agreement, while quieter or mismatched amplitudes led to disagreement. In the case of pitch, the significant contrasts were between ‘b’ and ‘c’, and between ‘c’ and ‘d’, suggesting that pitch plays a more nuanced role in participant perception. For force thresholds, the most notable finding was that ‘a’ responses differed significantly from all other choices, implying that participants who strongly agreed tended to apply greater force. Overall, these results suggest that both auditory intensity and tactile input (force) strongly influenced perceived congruence, while speaker identity (pain sound) has little perceptual impact.
To visualize the relationship between these features, we inspected a PCA projection of the standardized predictors (Figure 8B). The projection illustrates that data points for agreement (Choice ‘b’) cluster densely, whereas strong agreement (‘a’) and strong disagreement (‘d’) diverge along the primary components. This visual structure complements the ANOVA results (Table 2), which indicate that Force Threshold and Amplitude are the statistically strongest drivers of participant choice, while Pitch plays a significant but secondary role. Pain Sound (speaker identity) showed no significant influence, confirming that participants prioritized the intensity of the physical and auditory cues over the specific voice identity.
@@
As shown in Figure 8B, participant responses form overlapping but non-random clusters in PCA space. Notably, data points associated with choice ‘b’ (agreement) appear more densely centered near the origin, suggesting that moderate values of features are generally associated with perceptual agreement. In contrast, choice ‘a’ (strong agreement) and choice ‘d’ (strong disagreement) tend to appear at opposite poles along PC1. This spatial distribution supports the hypothesis that the combination of sound intensity and force magnitude contributes more to consistent perceptual judgments than sound identity alone.
These findings indicate that perceived congruence in this study was heavily influenced by the intensity of the interaction (Amplitude and Force). Unlike the expectation that Pitch would be the primary driver of urgency, our data indicates that participants prioritized Force and Amplitude when determining the realism of the simulation, with Pitch playing a secondary, nuanced role. Speaker identity (Pain Sound) exerted negligible effect. These insights offer practical implications for robopatient design: to elicit strong user agreement, the system must prioritize synchronizing sound amplitude with applied force, while pitch variation may act as a subtle, rather than dominant, cue.